OpenStudy (anonymous):

What is the value of the discriminant of the polynomial below? x2 - 9

OpenStudy (anonymous):

For \(ax^2+bx+c\), the discriminant is \(b^2-4ac\). So, you have a=1, b=0, c=-9. What is the discriminant?

OpenStudy (unklerhaukus):

\[\Delta=b^2-4ac\]\[=0^2-4\times1\times-9\]

OpenStudy (unklerhaukus):

thirty six